8. Solve x

x^2y′′− 2xy′+ 2y = x^4sin(4 log x)


1
Expert's answer
2022-03-17T06:56:34-0400

"x=e^u"

"a_0\lambda(\lambda-1)(\lambda-2)...(\lambda-n+1)...a_{n-2}\lambda(\lambda-1)+a_{n-1}\lambda+a_n=0"

"(\lambda-1)\lambda-2\lambda+2=0"

"\lambda^2-3\lambda+2=0"

"\lambda_1=2"

"\lambda_2=1"

"y=Ce^u+C_1e^{2u}"

"y_i=u^se^{au}(R_m(u)cos(bu)+T_m(u)sin(bu))"

s=0 if a+bi is not a root or s=k if it is.

a+bi=4i+4 s=0

"y_0=e^{4u}(Bsin(4u)+Acos(4u))"

"y'_0(u)=(4B-4A)e^{4u}sin(4u)+(4B+4A)e^{4u}cos(4u)"

"y''_0=32Be^{4u}cos(4u)-32Ae^{4u}sin(4u)"

"(-10B-20A)e^{4u}sin(4u)+(20B-10A)e^{4u}cos(4u)=e^{4u}sin(4u)"

"A=\frac{-1}{25}"

"B=\frac{-1}{50}"

"y_0=e^{4u}(\frac{-sin(4u)}{50}-\frac{cos(4u)}{25})"

"y=e^{4u}(\frac{-sin(4u)}{50}-\frac{cos(4u)}{25})+Ce^u+C_1e^{2u}"